OpenStudy (anonymous):

w(t)=3t-1. Find w(9)

OpenStudy (anonymous):

find domain of f(x)=x+3/x+2

OpenStudy (anonymous):

anybody know how to do this

OpenStudy (jdoe0001):

what's say 25/0?

OpenStudy (jdoe0001):

can you actually get a value from \(\bf \cfrac{25}{0}\) ?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

so it's undefined?

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

The domain can be ANY real number except -2 Because then it can be UNDEFINED! Understood? :)

OpenStudy (jdoe0001):

\(\bf f(x) = \cfrac{x+3}{x+2}\) if the denominator in a fraction becomes "0", then then the fraction becomes undefined, yes thus when taking that into account for the DOMAIN, or values "x" can take we have to keep in mind that the denominator of the rational mustn't be undefined so the values "x" can take are the ones that won't make the denominator 0

OpenStudy (anonymous):

so (-1,infintity)?

OpenStudy (jdoe0001):

look at AkashdeepDeb line above :)

OpenStudy (anonymous):

i get that but how do i say that

OpenStudy (jdoe0001):

if x= -2, then x+2 becomes (-2) + 2 = 0 if you give "x" any other value, negative or positive, it won't make the denominator 0

OpenStudy (anonymous):

ok. so my domain is all real numbers other than -2

OpenStudy (jdoe0001):

well, I guess I could put it as \(\bf \large x \in \mathbb{R}; x \ne -2\)

OpenStudy (anonymous):

ok thanks